# Information-Theoretic Thresholds for Bipartite Latent-Space Graphs Under Noisy Observations

### v1

@@ 1. Introduction
Latent geometric structure is a common feature in large-scale datasets appearing in various domains including data science [10], statistical physics [12] or biological sciences [7]. Random geometric graphs (RGGs) where vertices are represented by points randomly distributed in some latent geometric space and connected as a function of their distance, provide a natural way of modeling such high-dimensional datasets. When the latent geometric space incorporates a natural measure of dimensionality and suitable symmetries, RGGs have been shown to converge (in total variation distance) to Erdős–Rényi graphs [^1] In the Erdős–Rényi model edges are drawn independently with probability $p\in(0,1)$. as the dimension $d$ tends to infinity [8]. This prompts the fundamental question of finding the precise regimes of parameters $d,n,$ and edge-density $p$ in which RGGs are distinguishable from “pure noise” as modeled by Erdős–Rényi random graphs of the same edge-density, or as Duchemin and De Castro [9] state it: “determine the point where the geometry is lost in dimension”.
@@
In this direction, one of the most canonical and widely-studied models is that of Gaussian RGGs, which is obtained by associating a latent vector $\mathbf{x}_{v}$ to every vertex $v$ which is independently drawn from the standard $d$-dimensional Gaussian distribution $\mathcal{N}(0,\mathbf{I}_{d})$. Edges are inserted whenever the inner product $\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle$ exceeds (or undershoots) a certain threshold $\tau$. The resulting adjacency matrix can also be seen as a discretized Wishart matrix, and the model further shares strong similarities with spherical RGGs where vertices are represented by points on $\mathbb{S}^{d-1}$, i.e. the surface of the $d$-dimensional sphere, drawn from the Haar measure.
@@
In the dense regime, where the edge-density $p\in(0,1)$ is fixed, it is known that Gaussian RGGs become information-theoretically indistinguishable from Erdős–Rényi once $d\gg n^{3}$, while $d\ll n^{3}$ ensures existence of a simple, efficient test, formed by signed triangles  [6, 13, 19]. However, understanding variations of this problem that involve some kind of sparsity or additional noise have turned out to be much more challenging. In this direction, a series of recent works was concerned with the sparse case (where $p=o(1)$), and showed that the thresholds shift as a function of $p$ [15, 5, 3]. Nonetheless, obtaining tight information-theoretic thresholds for all $p$ is still a striking open problem. In a different direction, the works [17, 16, 4] have considered different notions of sparsity and noisy settings where $p$ is a fixed constant, while only a fraction of edges carries latent information, while the other entries are i.i.d. $\mathsf{Bern}(p)$ random variables [^2] This exact setting is studied in [17], while the work [4] studies a related question for wishart matrices, and [16] considers different notions of noise, parameterized by the “smoothness” of the connection function. However, like in the case of $p=o(1)$, finding the precise information-theoretic thresholds for all noise levels is still an open problem, as the bounds given in [17] leave some gaps.  [^3] The bounds given in [4] are tight, but the model is continuous and assumes that the “masked” edges are known up-front, which is in contrast to [17].
@@
In this work, we make some progress along the latter line of research, by giving tight information theoretic thresholds in an important special case of the model, formed by bipartite Gaussian RGGs. This constitutes progress towards resolving the gaps left open in [17], and further reveals striking differences between the setting where edges that carry latent information are revealed up-front as considered in [4], and the case where this information is concealed like in [17]. They further allow us to rule out the existence of computational statistical gaps. Moreover, we believe the techniques introduced here extend beyond the scope of this work, and yield a better understanding of the sparse case where $p=o(1)$.
@@ 1.1. The model and associated testing problems
We denote by $\mathbb{W}(n,m,p,d)$ the following distribution over $n\times m$ matrices with entries in $\{0,1\}$, where we refer to the set of rows as $R$ and to the set of columns as $L$, with $|R|=n$ and $|L|=m$. Without loss of generality, we further assume throughout that $m\geq n$.
@@
Moreover, we define $\mathbb{M}(n,m,p)$ as as the distribution over $n\times m$ matrices with entries in $\{0,1\}$ where each entry is an i.i.d. sample from the $\mathsf{Bern}(p)$ distribution.
Note that in $W\sim\mathbb{W}(n,m,p,d)$, each entry is marginally a $\mathsf{Bern}(p)$ random variable, and that the distribution $\mathbb{W}(n,m,p,d)$ can also be seen as the distribution over bipartite Gaussian RGGs with edge-density $p$. What is interesting about $W\sim\mathbb{W}(n,m,p,d)$ is that there are intricate dependencies are present within $W$ as all the entries depend on the same underlying latent information, whose strength crucially depends on the dimension $d$.
@@
As briefly mentioned earlier, previous works (see for example [6, 4]) have asked the natural question of which values of $d$ as a function of $n$, a matrix $M$ sampled from $\mathbb{W}(n,m,p,d)$ and closely related distributions can be distinguished from a typical sample $M\sim\mathbb{M}(n,m,p)$. In case of distinguishing $\mathbb{W}_{\mathbf{M}}(n,m,p,d)$ from $\mathbb{M}(n,m,p)$, it follows from the work of Brennan, Bresler and Huang [4, Corollary 4.3] that $\mathrm{d}_{\text{TV}}\left(\mathbb{W}(n,m,p,d),\mathbb{M}(n,m,p)\right)\rightarrow 0$ if $d\gg nm$ while $\mathrm{d}_{\text{TV}}\left(\mathbb{W}(n,m,p,d),\mathbb{M}(n,m,p)\right)\rightarrow 1$ if $d\ll nm$ and $p\in(0,1)$ is any fixed constant.
Introducing sparsity via random masks of density $q$. While this result fully characterizes the problem in terms of its feasibility as a function of $d$, it is much more challenging to find the right thresholds if the considered matrix additionally exhibits some form of sparsity in the sense that only a small fraction of entries actually carries information about the underlying latent vectors. One of the most natural ways of introducing sparsity to the above problem is the so called masked model, as introduced in [4]. Here, there is some partially hidden matrix $M$ and we are given as input a mask $\mathbf{M}\sim\mathbb{M}(n,m,q)$ together with the matrix $\mathbf{M}\odot M$ (where $\odot$ denotes the Hadamard product), while the goal is to distinguish the following two hypotheses.
@@
Accordingly, an algorithm for solving the above testing problem only has access to the entries of $M\sim\mathbb{M}(n,m,p)$ or $M\sim\mathbb{W}(n,m,p,d)$ that are not “hidden” by the mask $\mathbf{M}$ (i.e., correspond to $\mathbf{M}_{u,v}=1$). Accordingly, the masked model introduces a new parameter $q$ that controls the amount of sparsity and influences the associated thresholds for distinguishability. Tight information-theoretic thresholds for this masked setting as a function of $d$ and $q$ were given in [4] for a continuous version of the above problem, where a Wishart matrix is to be distinguished from a matrix with i.i.d. Gaussian entries.
Testing with known and unknown masks. The masked setting considered so far assumes that the mask $\bf{M}$ is given explicitly, which means that an algorithm has access to the information which entries in the input $M$ potentially carry latent information. It is very interesting to compare this setting to the case where the mask $\mathbf{M}$ is not known. To introduce these two settings formally, we define the following “masked” distributions.
@@
Given a matrix $\mathbf{M}_{u,v}\in\{0,1\}^{n\times m}$, define the distribution $\mathbb{W}_{\mathbf{M}}(n,m,p,d)$ as the distribution over matrixes $M\in\{0,1\}^{n\times m}$ obtained by sampling $W\sim\mathbb{W}(n,m,p,d)$, $B\sim\mathbb{M}(n,m,p)$ and then setting
@@
where $\odot$ denotes the Hadamard product. Moreover, define $\mathbb{W}(n,m,q,p,d)$ as the distribution $\mathbb{W}_{\mathbf{M}}(n,m,p,d)$ resulting from drawing $\mathbf{M}$ from $\mathbb{M}(n,m,q)$, i.e.,
@@
Accordingly, a sample $M\sim\mathbb{W}(n,m,q,p,d)$ is obtained by choosing a matrix $W\sim\mathbb{W}(n,m,p,d)$ and a mask $\mathbf{M}\sim\mathbb{M}(n,m,q)$, and subsequently setting $M_{u,v}$ corresponding to $\mathbf{M}_{u,v}=0$ to an independent sample from $\mathsf{Bern}(p)$. This second step serves the purpose of “hiding” the mask $\mathbf{M}$ by ensuring that every entry in $M$ has the same marginal distribution. This can equivalently be seen as a form of introducing noise, analogous to the setting considered in [17] for (non-bipartite) RGGs on the sphere.
@@
Note that above, the testing problem for known masks is introduced slightly differently than in the previous paragraph in the sense that the entries “hidden” by the mask are independent $\mathsf{Bern}(p)$ random variables in both hypotheses, while previously they were set to zero under both $H_{0}$ and $H_{1}$. However, it is not hard to see that from the lens of distinguishability, the two settings are equivalent, assuming that the mask is part of the input. Our definitions above have the advantage that they make it easier to compare the the two settings. We refer to Figure˜1 for an illustration of how the dependencies within a sample from $\mathbb{W}(n,m,q,p,d)$ behave and in particular become weaker as $d$ grows and $q$ tends to zero.
@@ 1.2. Results
The main result of this work is the establishment of precise, information-theoretic thresholds for the feasibility of ˜1.4 and ˜1.3 that are tight up to logarithmic factors.
@@
Our first main result concerns an almost tight characterization of information-theoretic distinguishability of $\mathbb{W}(n,m,q,p,d)$ and $\mathbb{M}(n,m,p)$ for all fixed $p$, and all dimensions $d=d(n)$ and noise levels $q=q(n)$. Writing $f\gg g$ for functions $f(n),g(n)$ whenever $\lim_{n\to\infty}f(n)/g(n)=\infty$, our first result reads as follows. Note that we assume $d\gg\log(n)^{3}$ throughout, simply because it simplifies some of our arguments. This assumption can likely be removed.
@@
Efficient tests and absence of computational-statistical gaps. The above upper and lower bounds on the total variation distance between the two matrix ensembles are almost tight and differ only by a factor of $\text{polylog}(m)$. In the regime where $\lim\mathrm{d}_{\text{TV}}\left(\textsc{W},\textsc{M}\right)=1$, there are two different regimes characterized by distinct, optimal tests. If $d\ll nmq^{4}$, then counting signed 4-cycles distinguishes $W$ and $M$, and if $d\ll m\sqrt{n}q^{2}$, then this holds for counting signed wedges. Formally, for a matrix $M\in\{0,1\}^{n\times m}$ these are given by
@@
respectively. Theorem˜1.5 shows that once we are in the regime where none of these statistics succeeds555or strictly speaking, if we exceed this regime by a factor of $\log(n)$, then no algorithm can distinguish the two distributions. Moreover, since $C_{4}(M)$ and $P_{2}(M)$ are clearly efficiently computable, this implies in particular that there are no computational statistical gaps in our problem.
The case $p=\frac{1}{2}$ vs. $p\neq\frac{1}{2}$. Moreover, Theorem˜1.5 highlights a somewhat surprising effect arising from the choice of $p$: For $p=\frac{1}{2}$, the regime in which $\lim\mathrm{d}_{\text{TV}}\left(W,M\right)=1$ is significantly smaller than for the case where $p$ is strictly larger or smaller than $\frac{1}{2}$. This is a consequence of the symmetry of the Gaussian distribution. In particular, if $p=\frac{1}{2}$, then $\tau=0$ which has the consequence that counting wedges has no statistical power anymore (even conditional on the latent vector $\mathbf{x}_{i}$) since the two entries $M_{i,k}$ and $M_{i,\ell}$ will be i.i.d. $\mathsf{Bern}(\frac{1}{2})$ random variables. Theorem˜1.5 shows that this effect is not specific to signed wedges, but that it is more inherent in the sense that no algorithm can distinguish $W$ and $M$ in the regime where signed four-cycles fail, even if signed wedges would succeed in the $p\neq\frac{1}{2}$ case. Hence, choosing $p=\frac{1}{2}$ makes the problem information-theoretically strictly harder compared to any other fixed choice of $p$.
Known vs. unknown masks. In the setting of Theorem˜1.5, we assume that the random mask determining which entries are to be re-randomized is not part of the input. We complement this assumption by a similar result concerning the setting where the mask is part of the input. The following theorem shows that the testing thresholds established in Theorem˜1.5 shift significantly under the assumption that the mask is known.
@@
If $p\neq\frac{1}{2}$ then:  $\begin{aligned} \hskip 22.76228pt\mathrm{d}_{\text{TV}}\left(\mathbb{W}_{\mathbf{M}}(n,m,p,d),\mathbb{M}(n,m,p)\right)&=\begin{cases}1-o(1)&\text{if }d\ll nmq^{2}\hskip 10.81218pt\text{ or }\hskip 2.84544ptd\ll m\sqrt{n}q\\ o(1)&\text{if }d\gg nmq^{4}\hskip 2.84544pt\text{ and }\hskip 2.84544ptd\gg m\sqrt{n}q^{2}\end{cases}\end{aligned}$
@@
Discrete vs. continuous models. Note that while the setting of known masks was already considered in [4] for the continuous analogue of our model, the above theorem shows that the discrete version of the problem considered here exhibits different thresholds once $q$ is sufficiently small, even if $p\neq\frac{1}{2}$. This is because in the continuous model, there are subtle differences in the marginal distribution of every entry $M_{u,v}$ when comparing $H_{0}$ and $H_{1}$. Testing for these differences allows us to distinguish the two models whenever $d\ll\sqrt{nmq}$. In our discrete version, the marginals match exactly and convergence in total variation therefore occurs earlier. More details on the differences of the discrete and the continuous setting can be found in Section˜2.2.
Convergence in total variation, but only for averaged masks. Abbreviating for now $\mathbb{W}=\mathbb{W}(n,m,q,p,d)$, $\mathbb{W}_{\textbf{M}}=\mathbb{W}_{\mathbf{M}}(n,m,p,d)$, and $\mathbb{M}=\mathbb{M}(n,m,p)$, then the bounds from Theorem˜1.5 and Theorem˜1.6 lead to the remarkable phenomenon that for all fixed $p$, there is a regime of the parameters $d,q$ where $\mathrm{d}_{\text{TV}}\left(\mathbb{W},\mathbb{M}\right)=o(1)$, while for the same $d$ and $q$, we have $\mathrm{d}_{\text{TV}}(\mathbb{W}_{\textbf{M}},\mathbb{M})=1-o(1)$ with high probability over the choice of $\mathbf{M}$. This means that for most fixed masks $\mathbf{M}$ of density $q$ (i.e. $\mathbf{M}\sim\mathbb{M}(n,m,q)$), the resulting conditional distribution $\mathbb{W}_{\textbf{M}}$ is efficiently distinguishable from $\mathbb{M}$, while in expectation over $\mathbf{M}$, convergence in total variation occurs. In particular, with high probability over $\mathbf{M}$, we have

### v3

@@ 1. Introduction
Latent geometric structure is a common feature in large-scale datasets appearing in various domains including data science [10], statistical physics [12] or biological sciences [7]. Random geometric graphs (RGGs) where vertices are represented by points randomly distributed in some latent geometric space and connected as a function of their distance, provide a natural way of modeling such high-dimensional datasets. When the latent geometric space incorporates a natural measure of dimensionality and suitable symmetries, RGGs have been shown to converge (in total variation distance) to Erdős–Rényi graphs [^1] In the Erdős–Rényi model edges are drawn independently with probability $p\in(0,1)$. as the dimension $d$ tends to infinity [8]. This prompts the fundamental question of finding the precise regimes of parameters $d,n,$ and edge-density $p$ in which RGGs are distinguishable from Erdős–Rényi random graphs of the same edge-density, or as Duchemin and De Castro [9] state it: “determine the point where the geometry is lost in dimension”.
@@
In this direction, one of the most canonical and widely-studied models is that of Gaussian RGGs, which is obtained by associating a latent vector $\mathbf{x}_{v}$ to every vertex $v$ which is independently drawn from the standard $d$-dimensionals Gaussian distribution $\mathcal{N}(0,\mathbf{I}_{d})$. Edges are inserted whenever the inner product $\langle\mathbf{x}_{u},\mathbf{x}_{v}\rangle$ exceeds (or undershoots) a certain threshold $\tau$. The resulting adjacency matrix can also be seen as a discretized Wishart matrix, and the model further shares strong similarities with spherical RGGs where vertices are represented by points on $\mathbb{S}^{d-1}$, i.e. the surface of the $d$-dimensional sphere, drawn from the Haar measure.
@@
In the dense regime, where the edge-density $p\in(0,1)$ is fixed, it is known that Gaussian RGGs become information-theoretically indistinguishable from Erdős–Rényi once $d\gg n^{3}$, while $d\ll n^{3}$ ensures existence of a simple, efficient test, formed by signed triangles  [6, 13, 19]. However, understanding variations of this problem that involve some kind of sparsity or additional noise have turned out to be much more challenging. In this direction, a series of recent works was concerned with the sparse case (where $p=o(1)$), and showed that the thresholds for detectability shift as a function of $p$ [15, 5, 3]. Nonetheless, obtaining tight information-theoretic thresholds for all $p$ is still a striking open problem. In a different direction, the works [17, 16, 4] have considered other notions of sparsity where $p$ is a fixed constant, while only a fraction of edges carries latent information and the other entries are i.i.d. $\mathsf{Bern}(p)$ random variables [^2] This exact setting is studied in [17], while the work [4] studies a related question for wishart matrices, and [16] considers different notions of noise, parameterized by the “smoothness” of the connection function. However, like in the case of $p=o(1)$, finding the precise information-theoretic thresholds for all levels of sparsity/noise is still an open problem, as the bounds given in [17] leave some gaps open  [^3] The bounds given in [4] are tight, but the model is continuous and assumes that the “masked” edges are known up-front, which is in contrast to [17]..
@@
In this work, we make progress along the latter line of research, by giving tight information theoretic thresholds in an important special case of the models considered in previous work, formed by bipartite Gaussian RGGs. In this setting, we close the gaps left open in [17], and our bounds further reveal striking differences between the setting where edges that carry latent information are revealed up-front (as considered in a similar form in [4], also for the bipartite case), and the case where this information is concealed, i.e., not accessible for the testing procedure (like in [17]). Our results further allow us to rule out the existence of computational statistical gaps for all ranges of parameters in our model. Moreover, we hope that the techniques introduced here extend beyond the scope of this work, and believe that they might also yield a better understanding of other variations of our detection problem, such as the case of $p=o(1)$.
@@ 1.1. The model and associated testing problems
We proceed by introducing the concrete testing problems we will be working with. To this end, denote by $\mathbb{W}(n,m,p,d)$ the following distribution over $n\times m$ matrices with entries in $\{0,1\}$, where we refer to the set of rows as $R$ and to the set of columns as $L$, with $|R|=n$ and $|L|=m$. Without loss of generality, we further assume throughout that $m\geq n$.
@@
[…]
Note that in $W\sim\mathbb{W}(n,m,p,d)$, each entry is marginally a $\mathsf{Bern}(p)$ random variable just like in $\mathbb{M}(n,m,p)$, however the difference is that there are intricate dependencies within $W\sim\mathbb{W}(n,m,p,d)$ as all the entries depend on the same underlying latent information, whose strength crucially depends on the dimension $d$. A matrix $W\sim\mathbb{W}(n,m,p,d)$ can also be seen as the adjacency matrix of a bipartite Gaussian RGGs with edge-density $p$, while $M\sim\mathbb{M}(n,m,p)$ is distributed like the adjacency matrix of a bipartite Erdős–Rényi random graph.
@@
As briefly mentioned earlier, previous works (see for example [6, 4]) have been concerned with understanding for which values of $d$ (as a function of $n$), a matrix $W$ sampled from $\mathbb{W}(n,m,p,d)$ and closely related distributions can be distinguished from a typical sample $M\sim\mathbb{M}(n,m,p)$. In this regard, it follows from the work of Brennan, Bresler and Huang [4, Corollary 4.3] that the total variation distance $\mathrm{d}_{\text{TV}}\left(\mathbb{W}(n,m,p,d),\mathbb{M}(n,m,p)\right)\rightarrow 0$ if $d\gg nm$ while $\mathrm{d}_{\text{TV}}\left(\mathbb{W}(n,m,p,d),\mathbb{M}(n,m,p)\right)\rightarrow 1$ if $d\ll nm$ and $p\in(0,1)$ is any fixed constant.
[…]
@@
Accordingly, any testing procedure can only observe the subset of entries in $M$ that is not “hidden” by the mask $\mathbf{M}$ (i.e. those entries that correspond to $\mathbf{M}_{u,v}=1$). Therefore, the masked model introduces a new parameter $q$ that controls the number of observable entries (or the sparsity of the resulting masked matrix $\mathbf{M}\odot M$) and heavily influences the resulting thresholds for distinguishability. Tight information-theoretic thresholds for testing in this masked setting as a function of $d$ and $q$ were given in [4] for a continuous version of the above problem, where a Wishart matrix is to be distinguished from a matrix with i.i.d. Gaussian entries.
Testing with known and unknown masks. The masked setting considered so far assumes that the mask $\bf{M}$ is given explicitly, which means that a testing procedure “knows” which entries in the input $M$ carry latent information and which do not. It is very natural to ask what happens if this is not the case, i.e., if the mask $\mathbf{M}$ is not given explicitly such that one cannot clearly separate the masked and non-masked edges a-priori. To introduce this new setting formally and compare it to the original masked model, we define the following two distributions.
@@
Given a matrix $\mathbf{M}\in\{0,1\}^{n\times m}$, define the distribution $\mathbb{W}_{\mathbf{M}}(n,m,p,d)$ as the distribution over matrices $M\in\{0,1\}^{n\times m}$ obtained by sampling $W\sim\mathbb{W}(n,m,p,d)$, $B\sim\mathbb{M}(n,m,p)$ and then setting
@@
where $\odot$ denotes the Hadamard product. Moreover, define $\mathbb{W}(n,m,q,p,d)$ as the distribution $\mathbb{W}_{\mathbf{M}}(n,m,p,d)$ resulting after drawing $\mathbf{M}$ from $\mathbb{M}(n,m,q)$, i.e.,
@@
Accordingly, a sample $M\sim\mathbb{W}(n,m,q,p,d)$ is obtained by choosing a matrix $W\sim\mathbb{W}(n,m,p,d)$ and a mask $\mathbf{M}\sim\mathbb{M}(n,m,q)$, and subsequently re-randomizing each entry $M_{u,v}$ corresponding to $\mathbf{M}_{u,v}=0$ by replacing it with an independent sample from $\mathsf{Bern}(p)$. This re-randomization step serves the purpose of “hiding” the mask $\mathbf{M}$ by ensuring that every entry in $M$ has the same marginal distribution. This can equivalently be seen as a form of introducing noise, and it reproduces the setting considered in [17] for (non-bipartite) RGGs on the sphere.
@@
Note that the above testing problem for known masks is introduced slightly differently than in the previous paragraph in the sense that the entries “hidden” by the mask are independent $\mathsf{Bern}(p)$ random variables in both hypotheses, while previously they were set to zero under both $H_{0}$ and $H_{1}$. However, it is not hard to see that from the lens of distinguishability, the two settings are equivalent, assuming that the mask is part of the input. Our way of stating ˜1.4 and ˜1.3 has the advantage that it allows us to to compare the the two settings more easily. To get an intuition for how the dependencies within a sample from $\mathbb{W}(n,m,q,p,d)$ behave and in particular become weaker as $d$ grows and $q$ tends to zero, we refer to Figure˜1.
@@ 1.2. Results
The main result of this work is the establishment of precise, information-theoretic thresholds for the feasibility of ˜1.4 and ˜1.3 that are tight up to logarithmic factors. Before stating them formally, we introduce some further notation. To this end, we remark that we write $f\gg g$ for functions $f(n),g(n)$ whenever $\lim_{n\to\infty}f(n)/g(n)=\infty$. Furthermore, we use standard Landau-notation to characterize the asymptotic behavior of functions, and we denote graphs by lower-case Greek letters. A graph $\alpha$ is represented as a set of edges such that $|\alpha|$ denotes the number of edges, while $V(\alpha)$ is the corresponding set of vertices and $|V(\alpha)|$ is its cardinality.
With this, our first result concerns the feasibility of ˜1.3 and reads as follows. Note that we assume $d\gg\log(n)^{3}$ throughout, simply because it simplifies some of our arguments, in particular related to the algorithmic upper bounds. However, we strongly believe that this assumption can be removed at the expense of a higher technical complexity.
@@
Efficient tests and absence of computational-statistical gaps. The above upper and lower bounds on the total variation distance between the two matrix ensembles are almost tight and differ only by a factor of $\log(n)$. In the regime where $\lim\mathrm{d}_{\text{TV}}\left(\mathbb{W},\mathbb{M}\right)=1$, there are two different regimes characterized by distinct, optimal tests. If $d\ll nmq^{4}$, then counting signed 4-cycles distinguishes $W$ and $M$, and if $d\ll m\sqrt{n}q^{2}$, then this holds for counting signed wedges. Formally, for a matrix $M\in\{0,1\}^{n\times m}$ these are given by
@@
respectively. Theorem˜1.5 shows that once we are in the regime where none of these statistics succeeds555or strictly speaking, if we exceed this regime by a factor of $\log(n)$, then no algorithm can distinguish the two distributions. Moreover, since $C_{4}(M)$ and $P_{2}(M)$ are clearly efficiently computable, this implies in particular that there are no computational statistical gaps.
The case $p=\frac{1}{2}$ vs. $p\neq\frac{1}{2}$. Moreover, Theorem˜1.5 highlights a somewhat surprising effect arising from the choice of $p$: for $p=\frac{1}{2}$, the regime in which $\lim\mathrm{d}_{\text{TV}}\left(\mathbb{W},\mathbb{M}\right)=1$ is significantly smaller than for the case where $p$ is strictly larger or smaller than $\frac{1}{2}$. This is a consequence of the symmetry of the Gaussian distribution. In particular, if $p=\frac{1}{2}$, then $\tau=0$ which implies that counting wedges has no statistical power since the two entries $M_{i,k}$ and $M_{i,\ell}$ will be i.i.d. $\mathsf{Bern}(\frac{1}{2})$ random variables, even conditional on a concrete latent vector $\mathbf{x}_{i}$. Theorem˜1.5 shows that this effect is not specific to signed wedges, but that it is inherent in the sense that no algorithm can distinguish $W$ and $M$ in the regime where signed four-cycles fail, even if signed wedges would succeed if $p\neq\frac{1}{2}$. Hence, choosing $p=\frac{1}{2}$ makes ˜1.3 strictly harder compared to any other fixed choice of $p$.
Known vs. unknown masks. In the setting of Theorem˜1.5, we assume that the random mask determining which entries are to be re-randomized is not part of the input. The following theorem shows that the testing thresholds established in Theorem˜1.5 shift significantly if this is not the case, i.e., if we instead consider ˜1.4.
@@
If $p\neq\frac{1}{2}$ then:  $\begin{aligned} \hskip 22.76228pt\mathrm{d}_{\text{TV}}\left(\mathbb{W}_{\mathbf{M}}(n,m,p,d),\mathbb{M}(n,m,p)\right)&=\begin{cases}1-o(1)&\text{if }d\ll nmq^{2}\hskip 36.98866pt\text{ or }\hskip 2.84544ptd\ll m\sqrt{n}q\\ o(1)&\text{if }d\gg nmq^{2}\log(n)\hskip 2.84544pt\text{ and }\hskip 2.84544ptd\gg m\sqrt{n}q\log(n)\end{cases}\end{aligned}$
@@
Discrete vs. continuous models. Note that while the setting of known masks was already considered in [4] for the continuous analogue of our model, the above theorem shows that the discrete version of the problem considered here exhibits different thresholds once $q$ is sufficiently small, even if $p\neq\frac{1}{2}$. This is because in the continuous model, there are subtle differences in the marginal distribution of every entry $M_{u,v}$ when comparing $H_{0}$ and $H_{1}$. Testing for these differences enables an efficient test whenever $d\ll\sqrt{nmq}$. In our discrete version, the marginals match exactly and convergence in total variation therefore occurs earlier. More details on the differences of the discrete and the continuous setting can be found in Section˜2.2.
Convergence in total variation, but only for averaged masks. Abbreviating for now $\mathbb{W}=\mathbb{W}(n,m,q,p,d)$, $\mathbb{W}_{\textbf{M}}=\mathbb{W}_{\mathbf{M}}(n,m,p,d)$, and $\mathbb{M}=\mathbb{M}(n,m,p)$, then the bounds from Theorem˜1.5 and Theorem˜1.6 lead to the remarkable phenomenon that for all fixed $p$, there is a regime of the parameters $d,q$ where $\mathrm{d}_{\text{TV}}\left(\mathbb{W},\mathbb{M}\right)=o(1)$, while for the same $d$ and $q$, we have $\mathrm{d}_{\text{TV}}(\mathbb{W}_{\textbf{M}},\mathbb{M})=1-o(1)$, with high probability over the choice of $\mathbf{M}$. This means that for most fixed masks $\mathbf{M}$ of density $q$ (i.e. $\mathbf{M}\sim\mathbb{M}(n,m,q)$), the resulting conditional distribution $\mathbb{W}_{\textbf{M}}$ is efficiently distinguishable from $\mathbb{M}$, while in expectation over $\mathbf{M}$, convergence in total variation occurs. In particular, with high probability over $\mathbf{M}$, we have